# Hierarchical Time Series Forecasting with Robust Reconciliation

### v1

@@ 4.2 Experimental setup
Hierarchical forecasting first requires a set of base forecasts (Base), which do not account for the hierarchical structure. We generated these base forecasts using Prophet (Taylor and Letham, 2018), an open-source time series library from Meta, with its default settings. As a preliminary experiment, we also performed base forecasts using other time series forecasting methods such as ARIMA and LightGBM, but Prophet showed the best results. Our comparative methods included the bottom-up (BU), top-down (TD), GLS reconciliation (GLS), and MinT reconciliation (MinT) approaches introduced in Section˜2.
For our proposed method (Robust), the optimization problem (6) was solved using the MOSEK solver (MOSEK ApS, 2025). Moreover, we needed to decide two parameters for bootstrap to design uncertainty set, which was the number of sampleing $N_{B}$ and a width of the uncertainty set $\alpha$. In this experiments, $N_{B}$ was fixed at $5000$ and $\alpha$ was determined by validation from among the candidates specified in advance. For validation, we divided the observation period data into train and validation data in a ratio of 9:1, and used the $\alpha$ with the smallest RMSE in the validation data. The candidates for $\alpha$ were $0.5$, $0.6$, $0.7$, $0.8$, $0.9$, and $1.0$.
@@
$\mathrm{MAE}=\frac{1}{T^{\prime}}\sum_{\tau=T+1}^{T+T^{\prime}}\left|y_{X}^{\left(\tau\right)}-\tilde{y}_{X}^{\left(\tau\right)}\right|,~\mathrm{RMSE}=\sqrt{\frac{1}{T^{\prime}}\sum_{\tau=T+1}^{T+T^{\prime}}\left(y_{X}^{\left(\tau\right)}-\tilde{y}_{X}^{\left(\tau\right)}\right)^{2}}$
@@ 4.3 Results and discussion
The experimental results for all datasets are summarized in Table˜58. To make the results easier to understand, we calculated the ratio of MAE and RMSE for each hierarchical time series forecsting method when the MAE and RMSE of base forecast are set to $1$, and then calculated the mean and standard deviation for series included in the same hierarchical level. We defined these metrics as relative MAE and relative RMSE, respectively. The original MAE and RMSE of all series are listed in the appendix. In the tables, the format is “mean $\pm$ standard deviation.” Since top level of hierarchical structure (Level 0) includes only one series, the standard deviation for the series is not shown. The underlined values indicate the best forecast accuracy in terms of the mean for the corresponding hierarchical level.
For the Australian births dataset, the validation to decide a parameter for bootstrap resulted in $\alpha=1.0$. The proposed method achieved the highest accuracy in most cases. In the top level, our proposed method was the only approach that achieved higer accuracy than base forecast. In the bottom level, considering the standard deviation, some series showed significant improvement in forecast accuracy with the proposed method.
@@
total $1.000$ $1.027$ $1.000$ $1.003$ $1.003$ $\underline{0.929}$
state $1.000$ $1.000\pm 0.000$ $1.518\pm 1.224$ $0.963\pm 0.104$ $0.985\pm 0.031$ $\underline{0.949\pm 0.146}$
@@
total $1.000$ $1.011$ $1.000$ $1.001$ $1.001$ $\underline{0.970}$
state $1.000$ $1.000\pm 0.000$ $1.300\pm 0.781$ $\underline{0.976\pm 0.069}$ $0.992\pm 0.019$ $\underline{0.976\pm 0.091}$
On the Australian tourism dataset, through the validation process, the bootstrap parameter was determined to be $\alpha=0.5$. Our proposed method yielded lower forecast accuracy than the existing GLS and MinT reconciliation methods except for the top level. This underperformance seemed to be due to the small discrepancy between the forecast error covariance matrix estimated from the observation period and the true covariance matrix in the forecast period. This reduced the benefit of explicitly accounting for covariance uncertainty, so our method was not the most effective for this type of data.
@@
total $1.000$ $2.315$ $1.000$ $1.125$ $0.663$ $\underline{0.647}$
state $1.000$ $1.277\pm 0.229$ $0.968\pm 0.371$ $0.635\pm 0.168$ $\underline{0.573\pm 0.125}$ $0.779\pm 0.415$
zone $1.000$ $1.000\pm 0.000$ $0.925\pm 0.373$ $0.623\pm 0.163$ $\underline{0.568\pm 0.155}$ $0.840\pm 0.396$
@@
total $1.000$ $2.132$ $1.000$ $1.105$ $0.698$ $\underline{0.638}$
state $1.000$ $1.227\pm 0.179$ $0.989\pm 0.360$ $0.680\pm 0.171$ $\underline{0.617\pm 0.125}$ $0.789\pm 0.377$
@@
For the Walmart sales, the validation procedure for selecting the bootstrap parameter yielded a value of $\alpha=0.9$. Our method achieved the highest forecast accuracy for all levels, with substantial performance gains over existing methods.
@@
total $1.000$ $1.141$ $1.000$ $1.012$ $0.964$ $\underline{0.884}$
state $1.000$ $1.132\pm 0.110$ $1.260\pm 0.800$ $1.000\pm 0.027$ $0.953\pm 0.045$ $\underline{0.864\pm 0.027}$
store $1.000$ $1.000\pm 0.000$ $1.601\pm 0.929$ $0.878\pm 0.071$ $0.859\pm 0.085$ $\underline{0.791\pm 0.145}$
@@
total $1.000$ $1.136$ $1.000$ $1.012$ $0.965$ $\underline{0.887}$
state $1.000$ $1.128\pm 0.110$ $1.237\pm 0.770$ $0.999\pm 0.027$ $0.952\pm 0.046$ $\underline{0.866\pm 0.030}$
store $1.000$ $1.000\pm 0.000$ $1.564\pm 0.877$ $0.880\pm 0.070$ $0.863\pm 0.081$ $\underline{0.798\pm 0.135}$
On the Swiss electricity demand datasets, as determined by validation, the bootstrap parameter was set to $\alpha=0.7$. Our proposed method achieved the highest forecast accuracy at the upper levels, but its accuracy deteriorated at the bottom level compared to the base forecast. In other words, in order to improve the forecast accuracy of the upper levels, the forecast accuracy of the bottom level was being sacrificed. Similar to the Australian tourism data, the gap between the true covariance matrix and the estimated covariance matrix was small, so the effect of robustification was considered to be small.
@@
grid $1.000$ $1.050$ $1.000$ $1.020$ $1.119$ $\underline{0.893}$
agg. $1.000$ $1.052\pm 0.038$ $0.990\pm 0.294$ $0.988\pm 0.009$ $1.004\pm 0.004$ $\underline{0.817\pm 0.187}$
sub-agg. $1.000$ $1.072\pm 0.066$ $1.393\pm 0.485$ $0.996\pm 0.013$ $1.016\pm 0.086$ $\underline{0.871\pm 0.200}$
meter $1.000$ $1.000\pm 0.000$ $1.809\pm 0.895$ $\underline{0.984\pm 0.102}$ $1.023\pm 0.171$ $1.039\pm 0.265$
@@
grid $1.000$ $1.025$ $1.000$ $1.011$ $1.052$ $\underline{0.948}$
agg. $1.000$ $1.020\pm 0.037$ $76\pm 0.148 0.990$$90\pm 0.003$ $1.012\pm 0.013$ $\underline{0.902\pm 0.115}$
sub-agg. $1.000$ $033\pm 0.055 1.1$$169\pm 0.312$ $0.992\pm 0.004$ $1.026\pm 0.063$ $\underline{0.931\pm 0.137}$
meter $1.000$ $1.000\pm 0.000$ $1.713\pm 0.757$ $\underline{0.978\pm 0.088}$ $1.020\pm 0.137$ $1.015\pm 0.217$
The numerical experiments on these four datasets demonstrated the effectiveness of our proposed method for hierarchical time series forecasting. Across most target series and hierarchical levels, the proposed approach consistently achieved higher forecast accuracy than existing techniques. However, its accuracy was slightly lower than that of the GLS and MinT reconciliation methods on a few datasets. As mentioned in the paragraph on the results of the Australian tourism data, when the gap between the true covariance matrix and the estimated covariance matrix is small, the proposed method is not considered superior. In fact, while validation yielded a large $\alpha$ for Australian births data and Walmart sales data where the proposed method was highly effective, it yielded a small $\alpha$ for Australian tourism data and Swiss electricity demand data where the method was less effective. Therefore, if you want to know in advance whether the proposed method is effective for the data, we recommend obtaining the parameter $\alpha$ from the observation period data and checking whether the $\alpha$ is large.
@@ 5 Conclusion
In this paper, we proposed a robust hierarchical time series forecasting method. This approach introduces an uncertainty set for the covariance matrix of forecast errors and minimizes the forecast error between the observation values and the coherent forecasts. Through numerical experiments, we demonstrated that the proposed method often provides more accurate forecasts than existing hierarchical time series forecasting methods, although its performance can vary across different datasets.
@@ Acknowledgments
This study was conducted as a part of the Data Analysis Competition hosted by the Joint Association Study Group of Management Science. The authors would like to thank the organizers and Research and Innovation Co., Ltd..
@@ Appendix A Full results of numerical experiments
total $3716.98$ $3818.08$ $3716.98$ $3726.75$ $3728.20$ $\underline{3453.87}$
ACT $79.19$ $79.19$ $\underline{23.98}$ $68.19$ $77.37$ $70.47$
AUS $1864.44$ $1864.44$ $1858.23$ $1854.67$ $1834.54$ $\underline{1727.37}$
NSW $430.46$ $430.46$ $937.69$ $422.76$ $414.65$ $\underline{376.57}$
NT $41.22$ $41.22$ $43.60$ $\underline{31.95}$ $39.70$ $36.66$
QLD $221.98$ $221.98$ $230.76$ $218.12$ $219.48$ $\underline{204.42}$
SA $106.12$ $106.12$ $328.78$ $97.96$ $100.94$ $\underline{85.08}$
TAS $\underline{40.78}$ $\underline{40.78}$ $163.65$ $48.17$ $43.43$ $54.79$
VIC $711.02$ $711.02$ $\underline{432.63}$ $701.14$ $702.36$ $677.66$
WA $483.61$ $483.61$ $\underline{179.59}$ $472.60$ $474.25$ $455.24$
@@
total $6433.27$ $6504.39$ $6433.27$ $6440.33$ $6441.38$ $\underline{6238.75}$
ACT $93.86$ $93.86$ $\underline{44.96}$ $84.08$ $92.21$ $86.18$
AUS $3221.01$ $3221.01$ $3216.54$ $3213.95$ $3199.22$ $\underline{3119.71}$
NSW $779.88$ $779.88$ $1168.37$ $773.80$ $767.51$ $\underline{739.04}$
NT $54.88$ $54.88$ $57.73$ $\underline{46.72}$ $53.51$ $50.69$
QLD $451.99$ $451.99$ $462.24$ $447.91$ $449.34$ $\underline{435.25}$
SA $165.37$ $165.37$ $354.97$ $158.39$ $160.96$ $\underline{147.93}$
TAS $\underline{55.31}$ $\underline{55.31}$ $169.64$ $61.64$ $57.54$ $67.60$
VIC $1215.67$ $1215.67$ $\underline{1005.35}$ $1208.26$ $1209.18$ $1191.31$
WA $630.09$ $630.09$ $\underline{383.28}$ $620.83$ $622.21$ $606.17$
@@
total $1507.73$ $3490.86$ $1507.73$ $1696.30$ $999.98$ $\underline{975.44}$
NSW $706.38$ $1056.55$ $\underline{336.42}$ $594.78$ $445.01$ $680.28$
NT $122.88$ $122.99$ $151.28$ $57.50$ $82.99$ $\underline{49.88}$
QLD $565.88$ $626.18$ $503.91$ $427.08$ $\underline{206.32}$ $210.33$
SA $213.29$ $278.99$ $90.94$ $82.64$ $83.73$ $\underline{82.31}$
TAS $132.18$ $132.50$ $158.77$ $\underline{67.01}$ $84.52$ $79.48$
VIC $509.36$ $828.63$ $532.14$ $416.27$ $\underline{347.01}$ $694.33$
WA $331.08$ $462.09$ $498.96$ $222.41$ $\underline{207.11}$ $450.89$
NSW_ACT $131.54$ $131.54$ $68.63$ $\underline{63.37}$ $68.05$ $70.95$
NSW_Nth $139.01$ $139.01$ $92.79$ $75.93$ $82.49$ $\underline{53.79}$
NSW_Sth $131.20$ $131.20$ $115.03$ $\underline{77.85}$ $94.87$ $168.27$
NSW_Metro $313.69$ $313.69$ $161.47$ $241.31$ $\underline{142.86}$ $285.20$
NSW_Nth_Coast $237.56$ $237.56$ $\underline{119.50}$ $160.60$ $130.42$ $179.47$
NSW_Sth_Coast $118.41$ $118.41$ $136.72$ $53.19$ $\underline{44.83}$ $119.64$
NT_Central $72.85$ $72.85$ $84.78$ $\underline{37.56}$ $51.82$ $43.39$
NT_Nth_Coast $50.14$ $50.14$ $67.35$ $\underline{30.65}$ $31.72$ $70.02$
QLD_Metro $360.78$ $360.78$ $209.51$ $318.68$ $197.90$ $\underline{176.87}$
QLD_Central_Coast $53.20$ $53.20$ $69.45$ $\underline{33.78}$ $38.48$ $36.41$
QLD_Inland $150.29$ $150.29$ $172.18$ $107.54$ $\underline{64.82}$ $78.66$
QLD_Nth_Coast $80.10$ $80.10$ $140.75$ $\underline{58.20}$ $63.47$ $113.08$
SA_Metro $127.55$ $127.55$ $\underline{34.11}$ $76.69$ $36.13$ $35.48$
SA_Inland $78.95$ $78.95$ $\underline{30.53}$ $34.33$ $32.78$ $32.01$
SA_West_Coast $27.14$ $27.14$ $24.60$ $31.01$ $\underline{16.42}$ $19.22$
SA_Sth_Coast $52.22$ $52.22$ $70.85$ $\underline{37.05}$ $40.02$ $77.66$
TAS_Nth_East $50.76$ $50.76$ $61.78$ $\underline{29.15}$ $32.34$ $34.77$
TAS_Sth $53.34$ $53.34$ $70.71$ $\underline{39.37}$ $41.03$ $47.04$
TAS_Nth_West $39.50$ $39.50$ $33.94$ $\underline{23.07}$ $25.23$ $61.00$
VIC_Nth_West $100.44$ $100.44$ $64.15$ $53.47$ $\underline{53.19}$ $65.82$
VIC_Nth_East $209.01$ $209.01$ $\underline{86.61}$ $127.61$ $98.84$ $107.58$
VIC_Metro $335.23$ $335.23$ $339.66$ $250.59$ $\underline{176.29}$ $352.00$
VIC_East_Coast $113.59$ $113.59$ $126.28$ $\underline{73.29}$ $79.66$ $157.49$
VIC_West_Coast $85.57$ $85.57$ $93.60$ $\underline{39.71}$ $43.26$ $47.45$
WA_West_Coast $237.14$ $237.14$ $326.91$ $\underline{171.87}$ $199.24$ $363.23$
WA_Sth $96.10$ $96.10$ $56.24$ $\underline{30.70}$ $30.76$ $62.60$
WA_Nth $130.13$ $130.13$ $115.82$ $50.24$ $\underline{35.39}$ $47.26$

### v2

@@ 4.2 Experimental setup
Hierarchical forecasting first requires a set of base forecasts (Base), which do not account for the hierarchical structure. We generated these base forecasts using Prophet (Taylor and Letham, 2018), an open-source time-series library from Meta, with its default settings. As a preliminary experiment, we also performed base forecasts using other time-series forecasting methods such as ARIMA and LightGBM, but Prophet showed the best results. Our comparative methods included the bottom-up (BU), top-down (TD), GLS reconciliation (GLS), and MinT reconciliation (MinT) approaches introduced in Section˜2.
For our proposed method (Robust), the optimization problem (8) was solved using the MOSEK solver (MOSEK ApS, 2025). Moreover, we needed to set two parameters for the bootstrap to design the uncertainty set, the number of sampling $N_{B}$ and the width of the uncertainty set $\alpha$. In these experiments, $N_{B}$ was fixed at $5000$ and $\alpha$ was determined by validation from among the candidates specified in advance. For validation, we divided the observation period data into train and validation data in a ratio of 9:1, and used the $\alpha$ with the smallest RMSE in the validation data. The candidates for $\alpha$ were $0.5$, $0.6$, $0.7$, $0.8$, $0.9$, and $1.0$.
@@
$\mathrm{MAE}=\frac{1}{T^{\prime}}\sum_{\tau=T+1}^{T+T^{\prime}}\left|y_{X}^{(\tau)}-\tilde{y}_{X}^{(\tau)}\right|,~\mathrm{RMSE}=\sqrt{\frac{1}{T^{\prime}}\sum_{\tau=T+1}^{T+T^{\prime}}\left(y_{X}^{(\tau)}-\tilde{y}_{X}^{(\tau)}\right)^{2}}$
@@ 4.3 Results and discussion
The experimental results for all datasets are summarized in Table˜58. To make the results easier to understand, we calculated the ratio of MAE and RMSE for each hierarchical time-series forecasting method when the MAE and RMSE of the base forecasts are set to $1$, and then calculated the mean and standard deviation for series included in the same hierarchical level. We defined these metrics as relative MAE and relative RMSE, respectively. The original MAE and RMSE of all series are listed in the appendix. In the tables, the format is “mean $\pm$ standard deviation.” Since the top level of hierarchical structure (Level 0) includes only one series, the standard deviation for the series is not shown. The underlined values indicate the best forecast accuracy in terms of the mean for the corresponding hierarchical level.
For the Australian births dataset, the validation to decide a parameter for bootstrap resulted in $\alpha=1.0$. The proposed method achieved the highest accuracy in most cases. In the top level, our proposed method was the only approach that achieved higher accuracy than the base forecast. In the bottom level, considering the standard deviation, some series showed significant improvement in forecast accuracy with the proposed method.
@@
total $1.000$ $1.034$ $1.000$ $1.003$ $0.999$ $\underline{0.923}$
state $1.000$ $1.000\pm 0.000$ $1.513\pm 1.302$ $0.955\pm 0.136$ $0.978\pm 0.043$ $\underline{0.944\pm 0.167}$
@@
total $1.000$ $1.014$ $1.000$ $1.001$ $1.000$ $\underline{0.968}$
state $1.000$ $1.000\pm 0.000$ $1.297\pm 0.831$ $\underline{0.970\pm 0.090}$ $0.989\pm 0.026$ $0.973\pm 0.103$
In the Australian tourism dataset, the bootstrap parameter was determined via validation to be $\alpha=0.5$. Our proposed method yielded lower forecast accuracy than the existing GLS and MinT reconciliation methods, except for the top level. This underperformance seems to be due to a slight discrepancy between the estimated covariance matrix and the true one in the forecast period, which reduced the benefit of explicitly accounting for covariance uncertainty.
@@
total $1.000$ $2.315$ $1.000$ $1.125$ $0.663$ $\underline{0.518}$
state $1.000$ $1.277\pm 0.247$ $0.968\pm 0.400$ $0.635\pm 0.181$ $\underline{0.573\pm 0.135}$ $0.793\pm 0.460$
zone $1.000$ $1.000\pm 0.000$ $0.925\pm 0.380$ $0.623\pm 0.166$ $\underline{0.568\pm 0.158}$ $0.843\pm 0.387$
@@
total $1.000$ $2.132$ $1.000$ $1.105$ $0.698$ $\underline{0.519}$
state $1.000$ $1.227\pm 0.193$ $0.989\pm 0.388$ $0.680\pm 0.185$ $\underline{0.617\pm 0.135}$ $0.810\pm 0.428$
@@
For the Walmart sales, the validation procedure for selecting the bootstrap parameter yielded $\alpha=0.9$. Our method achieved the highest forecast accuracy for all levels, with substantial performance gains over existing methods.
@@
total $1.000$ $1.135$ $1.000$ $1.011$ $0.966$ $\underline{0.810}$
state $1.000$ $1.123\pm 0.117$ $1.263\pm 0.986$ $0.999\pm 0.029$ $0.956\pm 0.049$ $\underline{0.801\pm 0.036}$
store $1.000$ $1.000\pm 0.000$ $1.618\pm 1.001$ $0.882\pm 0.069$ $0.868\pm 0.076$ $\underline{0.741\pm 0.117}$
@@
total $1.000$ $1.129$ $1.000$ $1.011$ $0.967$ $\underline{0.817}$
state $1.000$ $1.118\pm 0.116$ $1.241\pm 0.948$ $0.998\pm 0.030$ $0.955\pm 0.050$ $\underline{0.807\pm 0.038}$
store $1.000$ $1.000\pm 0.000$ $1.580\pm 0.945$ $0.885\pm 0.067$ $0.872\pm 0.070$ $\underline{0.751\pm 0.111}$
On the Swiss electricity demand datasets, as determined by validation, the bootstrap parameter was set to $\alpha=0.7$. Our proposed method achieved the highest forecast accuracy at the upper levels, but the accuracy deteriorated at the bottom level compared to the base forecast. In other words, the improvement at the upper levels came at the expense of accuracy at the bottom level. As with the Australian tourism data, the discrepancy between the true and estimated covariance matrices may have been small, limiting the overall impact of robustification.
@@
grid $1.000$ $1.041$ $1.000$ $1.022$ $1.110$ $\underline{0.888}$
agg. $1.000$ $1.045\pm 0.034$ $0.983\pm 0.422$ $0.983\pm 0.020$ $0.995\pm 0.017$ $\underline{0.808\pm 0.277}$
sub-agg. $1.000$ $1.072\pm 0.071$ $1.390\pm 0.549$ $0.999\pm 0.010$ $1.014\pm 0.099$ $\underline{0.867\pm 0.232}$
meter $1.000$ $1.000\pm 0.000$ $1.810\pm 0.921$ $\underline{0.987\pm 0.102}$ $1.023\pm 0.177$ $1.036\pm 0.275$
@@
grid $1.000$ $1.022$ $1.000$ $1.012$ $1.050$ $\underline{0.946}$
agg. $1.000$ $1.013\pm 0.045$ $0.970\pm 0.213$ $0.986\pm 0.008$ $1.005\pm 0.023$ $\underline{0.895\pm 0.171}$
sub-agg. $1.000$ $032\pm 0.058 1.1$$170\pm 0.359$ $0.995\pm 0.003$ $1.025\pm 0.071$ $\underline{0.929\pm 0.158}$
meter $1.000$ $1.000\pm 0.000$ $1.717\pm 0.781$ $\underline{0.981\pm 0.086}$ $1.021\pm 0.141$ $1.013\pm 0.227$
The numerical experiments on these four datasets demonstrated the effectiveness of our proposed hierarchical time-series forecasting method. Across most target series and hierarchical levels, the proposed approach achieved higher forecast accuracy than existing techniques. However, in a few cases, the proposed method’s prediction performance was slightly inferior to that of the GLS and MinT reconciliation methods.
As discussed in the Australian tourism data results, when the discrepancy between the estimated and true covariance matrices is small, incorporating the uncertainty set may lead to overly conservative forecasts, resulting in lower predictive performance. This is further supported by the parameter selection during validation: a larger $\alpha$ was chosen for the Australian births and Walmart sales datasets, where our method was effective. Conversely, a smaller $\alpha$ was selected for the Australian tourism and Swiss electricity demand datasets, where the method was less effective. This suggests that for these latter datasets, the estimated covariance matrix closely approximated the true covariance matrix, reducing the need to account for its uncertainty in the reconciliation process.
@@ 5 Conclusion
In this paper, we proposed a robust hierarchical time-series forecasting method. This approach introduces an uncertainty set for the inverse of covariance matrix of forecast errors and minimizes the forecast error between the observation values and the coherent forecasts. Through numerical experiments, we demonstrated that the proposed method often provides more accurate forecasts than existing hierarchical time-series forecasting methods, although its performance can vary across different datasets.
@@ Acknowledgments
This study was conducted as a part of the Data Analysis Competition hosted by the Joint Association Study Group of Management Science. The authors would like to thank the organizers and Research and Innovation Co., Ltd.
@@ Appendix A Full results of numerical experiments
total $3697.93$ $3822.66$ $3697.93$ $3710.04$ $3695.48$ $\underline{3413.01}$
ACT $79.34$ $79.34$ $\underline{23.95}$ $65.71$ $77.18$ $71.65$
AUS $1863.94$ $1863.94$ $1848.74$ $1851.83$ $1819.48$ $\underline{1706.96}$
NSW $433.36$ $433.36$ $934.11$ $423.74$ $410.68$ $\underline{372.41}$
NT $41.86$ $41.86$ $43.48$ $\underline{30.60}$ $39.88$ $36.31$
QLD $222.39$ $222.39$ $231.03$ $217.53$ $217.92$ $\underline{202.75}$
SA $107.33$ $107.33$ $328.02$ $97.16$ $100.02$ $\underline{83.49}$
TAS $\underline{40.31}$ $\underline{40.31}$ $163.39$ $49.53$ $43.70$ $55.08$
VIC $711.71$ $711.71$ $\underline{430.98}$ $699.45$ $700.14$ $675.65$
WA $479.92$ $479.92$ $\underline{180.03}$ $466.29$ $468.86$ $451.56$
@@
total $6419.88$ $6507.75$ $6419.88$ $6428.59$ $6418.09$ $\underline{6213.25}$
ACT $93.99$ $93.99$ $\underline{44.91}$ $81.93$ $92.04$ $87.08$
AUS $3220.69$ $3220.69$ $3209.85$ $3211.97$ $3188.33$ $\underline{3106.97}$
NSW $782.22$ $782.22$ $1165.31$ $774.65$ $764.60$ $\underline{736.15}$
NT $55.47$ $55.47$ $57.62$ $\underline{45.46}$ $53.67$ $50.50$
QLD $452.41$ $452.41$ $462.21$ $447.36$ $447.76$ $\underline{433.83}$
SA $166.40$ $166.40$ $354.25$ $157.74$ $160.18$ $\underline{146.77}$
TAS $\underline{54.89}$ $\underline{54.89}$ $169.39$ $62.80$ $57.77$ $67.82$
VIC $1216.24$ $1216.24$ $\underline{1004.19}$ $1207.09$ $1207.60$ $1190.46$
WA $626.98$ $626.98$ $\underline{383.23}$ $615.56$ $617.71$ $603.22$
@@
total $1507.73$ $3490.86$ $1507.73$ $1696.30$ $999.98$ $\underline{780.57}$
NSW $706.38$ $1056.55$ $\underline{336.42}$ $594.78$ $445.01$ $678.46$
NT $122.88$ $122.99$ $151.28$ $\underline{57.50}$ $82.99$ $71.12$
QLD $565.88$ $626.18$ $503.91$ $427.08$ $\underline{206.32}$ $219.97$
SA $213.29$ $278.99$ $90.94$ $82.64$ $83.73$ $\underline{78.47}$
TAS $132.18$ $132.50$ $158.77$ $67.01$ $84.52$ $\underline{62.70}$
VIC $509.36$ $828.63$ $532.14$ $416.27$ $\underline{347.01}$ $640.93$
WA $331.08$ $462.09$ $498.96$ $222.41$ $\underline{207.11}$ $503.88$
NSW_ACT $131.54$ $131.54$ $68.63$ $\underline{63.37}$ $68.05$ $102.13$
NSW_Nth $139.01$ $139.01$ $92.79$ $75.93$ $82.49$ $\underline{57.30}$
NSW_Sth $131.20$ $131.20$ $115.03$ $\underline{77.85}$ $94.87$ $165.99$
NSW_Metro $313.69$ $313.69$ $161.47$ $241.31$ $\underline{142.86}$ $323.77$
NSW_Nth_Coast $237.56$ $237.56$ $\underline{119.50}$ $160.60$ $130.42$ $190.16$
NSW_Sth_Coast $118.41$ $118.41$ $136.72$ $53.19$ $\underline{44.83}$ $110.12$
NT_Central $72.85$ $72.85$ $84.78$ $37.56$ $51.82$ $\underline{32.97}$
NT_Nth_Coast $50.14$ $50.14$ $67.35$ $\underline{30.65}$ $31.72$ $76.22$
QLD_Metro $360.78$ $360.78$ $209.51$ $318.68$ $197.90$ $\underline{173.06}$
QLD_Central_Coast $53.20$ $53.20$ $69.45$ $\underline{33.78}$ $38.48$ $60.19$
QLD_Inland $150.29$ $150.29$ $172.18$ $107.54$ $\underline{64.82}$ $84.05$
QLD_Nth_Coast $80.10$ $80.10$ $140.75$ $\underline{58.20}$ $63.47$ $128.42$
SA_Metro $127.55$ $127.55$ $\underline{34.11}$ $76.69$ $36.13$ $38.15$
SA_Inland $78.95$ $78.95$ $\underline{30.53}$ $34.33$ $32.78$ $38.62$
SA_West_Coast $27.14$ $27.14$ $24.60$ $31.01$ $\underline{16.42}$ $22.17$
SA_Sth_Coast $52.22$ $52.22$ $70.85$ $\underline{37.05}$ $40.02$ $57.76$
TAS_Nth_East $50.76$ $50.76$ $61.78$ $\underline{29.15}$ $32.34$ $39.28$
TAS_Sth $53.34$ $53.34$ $70.71$ $\underline{39.37}$ $41.03$ $46.34$
TAS_Nth_West $39.50$ $39.50$ $33.94$ $23.07$ $25.23$ $\underline{21.32}$
VIC_Nth_West $100.44$ $100.44$ $64.15$ $53.47$ $\underline{53.19}$ $59.12$
VIC_Nth_East $209.01$ $209.01$ $\underline{86.61}$ $127.61$ $98.84$ $87.13$
VIC_Metro $335.23$ $335.23$ $339.66$ $250.59$ $\underline{176.29}$ $335.11$
VIC_East_Coast $113.59$ $113.59$ $126.28$ $\underline{73.29}$ $79.66$ $154.80$
VIC_West_Coast $85.57$ $85.57$ $93.60$ $\underline{39.71}$ $43.26$ $59.43$
WA_West_Coast $237.14$ $237.14$ $326.91$ $\underline{171.87}$ $199.24$ $392.71$
WA_Sth $96.10$ $96.10$ $56.24$ $\underline{30.70}$ $30.76$ $70.90$
WA_Nth $130.13$ $130.13$ $115.82$ $50.24$ $\underline{35.39}$ $57.09$